# Moyun: A Diffusion-Based Model for Style-Specific Chinese Calligraphy Generation

### v1

@@
Abstract
Although Chinese calligraphy generation has achieved style transfer, generating calligraphy by specifying the calligrapher, font, and character style remains challenging. To address this, we propose a new Chinese calligraphy generation model “Moyun” , which replaces the Unet in the Diffusion model with Vision Mamba and introduces the TripleLabel control mechanism to achieve controllable calligraphy generation. The model was tested on our large-scale dataset “Mobao” of over 1.9 million images, and the results demonstrate that “Moyun” can effectively control the generation process and produce calligraphy in the specified style. Even for calligraphy the calligrapher has not written, “Moyun” can generate calligraphy that matches the style of the calligrapher.
Index Terms:
Calligraphy, Diffusion, Mamba
I Introduction
Chinese calligraphy, with a history spanning over thousands of years, is a cherished cultural treasure of China which represents the artistic of Chinese characters handwriting. Chinese calligraphy is rich in variations. Chinese has tens of thousands of characters, each with a different meaning. Additionally, a single character can be written in various fonts, such as regular script, running script, cursive script, clerical script, seal script, and so on. Moreover, the writing of the same font differs between calligraphers, as shown in Figure 1a. The writing of the same font by the same calligrapher shows consistency, which we refer to as calligraphic style. The rich variations in Chinese calligraphy require an extended period of study for an average person to master. However, Chinese calligraphy is widely used, which has led people to explore the use of AI for generating Chinese calligraphy. Recently, GAN and Diffusion models are applied to Chinese calligraphy generation [1, 2, 3, 4], yielding impressive outcomes.
Figure 1: (a) shows the character “bai” (which means ”white” in Chinese) written in different fonts by various calligraphers. Each column represents a different font, and each row corresponds to a different calligrapher. (b) The first row shows calligraphy generated by Calliffusion, while the second row shows the ground truth. In the first column (san, regular script, Yan Zhenqing), the strokes in the ground truth are evenly spaced, but Calliffusion’s result is not. In the second column (ba, cursive script, Su Shi), Calliffusion’s output appears less stable than the ground truth from an aesthetic standpoint.
@@ I Introduction
ZiGAN [2] is a seminal work that applies GAN to the field of Chinese calligraphy generation. ZiGAN constructed several small-scale datasets, each with consistent styles, and trained models separately on these datasets. As a result, the trained models gained the ability to transform standard printed characters into calligraphy [^1] In the following text, we use “Calligraphy” to refer to a single character image generated by the calligraphy generation model. with the style of the specific dataset used during training. However, it cannot generate calligraphy guided by specific calligraphers, fonts, and characters. Calliffusion [3] is the first Chinese calligraphy generation model based on diffusion model[5]. Calliffusion uses descriptive text involving “character, font, and calligrapher” to effectively guide the generation process. However, it relies on a BERT[6] model pre-trained on Chinese language dataset as the text encoder, which struggles to accurately understand domain-specific terms in calligraphy, such as the names of calligraphers. For example, a calligrapher’s name might be split into two tokens. Additionally, this approach introduces extra computational cost, which is unnecessary for calligraphy generation tasks. To better accomplish this task, we proposed a multilabel mechanism, where independent classification labels correspond to the calligrapher, font, and character, and these labels are combined to control the generation process.
Calliffusion uses an Diffusion architecture based on Unet[7], but Unet does not adequately fit the structural relationships between the strokes of the characters, which is shown in Figure 1b. Our model is based on the Vision Mamba[8], which processes images through patchify. Additionally, we incorporated the more efficient Mamba model[9]. Experiments demonstrate that our model provides a better fit to the structure between strokes.
The ZiGAN’s dataset[2] is relatively small, containing only 9 sub-datasets, with each sub-dataset consisting of 6,000 images of a single style. In contrast, we collected 1.9 million single-character calligraphy images with diverse styles and detailed annotations. Due to varying collection conditions and different noise distributions, traditional binarization methods performed poorly. To address this, we designed a new binarization method based on SAM[10], which achieved better results.
@@
II Related works
II-A Generative Adversarial Networks
Generative Adversarial Networks (GAN) [11] consists of a generator and a discriminator. The generator tries to produce data similar to real data, while the discriminator’s task is to distinguish between the generated data and real data. Research on GAN-based calligraphy generation has been ongoing for a long time, with the earliest work being zi2zi [4], an open-source project based on pix2pix [12]. After that, there were calliGAN [1], ZiGAN[2], and end-to-end[13] models. The most recent work is a style transfer model called CCST-GAN[14].
II-B Diffusion Model
The Diffusion model is a widely studied image generation model, with its structure initially established in DDPM [5] and DDIM [15], and the introduction of VAE[16] in LDM [17] further improving its application. DiT [18] brought the Transformer [19] architecture into the Diffusion model. Research on Diffusion-based calligraphy generation began relatively late. The earliest work is Calliffusion[3], which utilized a Unet-based Diffusion model. The most recent work is DP-Font[20], which employs a physical information neural network structure, achieving better control over character shapes.
III Method
@@
Figure 2 shows the architecture of “Moyun”. “Moyun” is a diffusion model based on Vision Mamba[8], optimized with Mamba2[9]. Furthermore, we control the generation process by the TripleLabel mechanism.
@@
Figure 2: “Moyun” architecture. The input latent noise is patched. The label is a combination of the calligrapher, font, and character. We used Mamba2-Replacement-Vision Mamba to process the patches.
III-A Preliminaries
@@
Moyun is a diffusion model, so our training and inference follow the diffusion [5, 15] approach. The forward process gradually adds noise to the real data, specifically to our calligraphy image $x_{0}$. The diffusion process follows the equation as shown in (1) .
$\vspace{-5pt}q(x_{t}|x_{0})=\mathcal{N}\left(x_{t};\bar{\alpha}_{t}x_{0},(1-\bar{\alpha}_{t})\mathbf{I}\right)\vspace{-5pt}$ (1)
@@
$x_{t}=\bar{\alpha}_{t}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\epsilon_{t}$ (2)
The diffusion model trains the reverse denoising process, and the denoising equation is as follows:
$p_{\theta}(x_{t-1}\mid x_{t})=\mathcal{N}(\mu_{\theta}(x_{t}),\Sigma_{\theta}(x_{t}))$ (3)
@@
$\mathcal{L}_{\text{simple}}(\theta)=\|\epsilon_{\theta}(x_{t})-\epsilon_{t}\|^{2}\vspace{-5pt}$ (4)
III-B Block Design
To improve the model’s ability to capture the structure of calligraphy, we replaced the Unet in the Diffusion Model with Vision Mamba[8], thereby introducing the patchify mechanism. Additionally, we used the latest Mamba2[9] model instead of the first version of Mamba[21] in Vision Mamba. Thus, each patch corresponds to a small part of the calligraphic structure, and Mamba’s efficient contextual relational ability effectively matches the relationships between patches, thereby better learning the structure of the calligraphy.
The input image is first mapped into the latent space via a VAE[16], followed by patchify processing. For a square image with dimensions of $a\times a$, it is first mapped to a latent space of size $b\times b$ by the variational autoencoder (VAE). The latent representation is then divided into $n$ patches of size $p\times p$, where $n=\frac{b^{2}}{p^{2}}$, with $p$ being the side length of each patch and $b$ the side length of the latent space. In total, the entire Block is iterated $N$ times.
III-C TripleLabel control
To accommodate the triple-conditions control requirements of Chinese calligraphy generation, which include calligrapher, font, and character, we designed a TripleLabel control mechanism. Each calligrapher, font, and character is mapped to a unique class label, represented by a number, and each label is arbitrary. This allows the combination of labels to generate calligraphy that the calligrapher has not written before. In the model, the input label is transformed into the corresponding embedding vector through an embedding table. The three resulting embedding vectors are summed and used to control the generation process via a scale-shift mechanism[18].
@@ III-C TripleLabel control
This control mechanism significantly reduces computational cost compared to introducing a new text encoder. Furthermore, subsequent experiments demonstrated that this method of control is highly effective.
IV experiment
IV-A Dataset
@@
Figure 3: (a) shows the directory structure of dataset “Mobao”, using the calligrapher “Fan Zhongyan” as an example. “Zu” and “Zi” only have single images, while “Bai” has multiple images. (b) demonstrates the binarization process, using the character “Zu” as an example to show the steps of selecting points, obtaining the mask, and resizing.
Dataset construct
To construct the large-scale annotated dataset “Mobao”, we scraped images from web and subject them to binarization processing. The collected calligraphy images include six fonts: regular script, running script, cursive script, clerical script, seal script, and seal carving. Each calligraphy image is annotated with the calligrapher, font, and character. We organized the data into a hierarchical folder structure of calligrapher-font-character, and each image within these categories was numbered. The complete processing workflow is illustrated in Figure 3.
To perform binarization, we employ SAM for image segmentation. Initially, Otsu binarization is applied to the image, and connected regions larger than 100 pixels are selected, which typically represent strokes in calligraphy. Within these connected regions, positive points are uniformly sampled using k-means. In addition, negative points are selected outside of these connected domains. The formula for selecting the number of points is as follows:
$\vspace{-5pt}cnt_{pos}=\max(1,\min(20,\frac{area_{regions}\times 100}{area_{total}})\vspace{-1pt}$ (5)
$\vspace{-1pt}cnt_{neg}=50$ (6)
These points are then provided to SAM as prompts with original calligraphy images, leading to a mask output by SAM, which we regard as the foreground. By filling the complement with white, we obtain the binarized image. Finally, a resizing process is conducted, in which the longer side of the image’s dimensions is stretched to 256 while maintaining the aspect ratio. Afterwards, it is centered and concatenated with a white background to yield a $256\times 256$ image.
dataset analyse
@@
After processing the dataset, we conducted an analysis. We obtained a total of 1,929,393 images, including works from 6 fonts, 2,681 calligraphers and 4,660 characters. Some images without attributed calligraphers were categorized under the “anonymous” category.
@@
To further evaluate our dataset, we conducted statistical analysis across the three dimensions of calligrapher, font, and character. The results confirm that our dataset is vast: the calligrapher with the most collected works, “Wang Xizhi”, has a total of 124,854 images; the font with the largest collection , running script, contains 668,040 images; and the character with the most instances, “shu” (book in chinese), appears in 11,949 images. This provides a rich source of data for the model to learn the structure and style of calligraphy.
However, our dataset has some issues. Nearly half of the calligraphers have fewer than ten collected works, and half of the characters have fewer than one hundred images. This indicates that our dataset exhibits a long-tail distribution.
IV-B Experiment Setup
Experiment Dataset
To ensure the accuracy of the experiment, we selected a balanced subset from the complete dataset to serve as the experimental dataset. The specific method is as follows: We chose 40 calligraphers and selected 40 characters that all of them have written. For each calligrapher, 90% of the characters were used for training, and the remaining characters were used for testing. This ensures that each test character is unseen during training for that particular calligrapher, while other characters by the same calligrapher and the same characters written by other calligraphers are used for training. This approach ensures that both the character shapes and the calligrapher’s style are adequately trained. Additionally, due to the varying number of characters for each calligrapher, we randomly selected up to four images for each character. Finally, we chose 12,985 images, including 11,689 images in tranining set and 1,296 images in test set. In this way, the negative impact of the long tail effect was successfully avoid .
Model Specifics

### v2

@@
Abstract.
Although Chinese calligraphy generation has made progress in style transfer or font synthesis, generating calligraphy by precisely specifying the calligrapher, font, and character style remains challenging. To address this, we propose a new Chinese calligraphy generation model “Moyun” , which replaces the UNet in the Diffusion model with Vision Mamba and introduces the TripleLabel control mechanism to achieve controllable calligraphy generation. The model was tested on our large-scale dataset “Mobao”, which contains over 1.9 million images. The results demonstrate that “Moyun” can effectively control the generation process and produce calligraphy in the specified style. Even for calligraphy the calligrapher has not written, “Moyun” can generate calligraphy that matches the style of the calligrapher.
Chinese Calligraphy, Diffusion Model, Mamba
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 3rd International Workshop on Multimedia Content Generation and Evaluation: New Methods and Practice; October 27–28, 2025; Dublin, Irelandbooktitle: Proceedings of the 3rd International Workshop on Multimedia Content Generation and Evaluation: New Methods and Practice (McGE ’25), October 27–28, 2025, Dublin, Irelanddoi: 10.1145/3746278.3759378isbn: 979-8-4007-2060-4/2025/10ccs: Computing methodologies Artificial intelligence
1. Introduction
Chinese calligraphy, with a history spanning over thousands of years, is a cherished cultural treasure of China which represents the artistic of Chinese characters handwriting. As a world intangible cultural heritage, its unique aesthetic characteristics are reflected in the lightness and heaviness of the brushstrokes, the thickness and dryness of the ink levels, and the virtual and real echoes of the layout. Chinese calligraphy is rich in variations. Chinese has tens of thousands of characters, each with a different meaning. Additionally, a single character can be written in various fonts, such as regular script, running script, cursive script, clerical script, seal script, and so on. Moreover, the writing of the same font differs between calligraphers, as shown in Figure 1a. The writing of the same font by the same calligrapher shows consistency, which we refer to as calligraphic style. The rich variations in Chinese calligraphy require an extended period of study for an average person to master. However, Chinese calligraphy is widely used, which has led people to explore the use of AI for generating Chinese calligraphy. Recently, GAN and Diffusion models are applied to Chinese calligraphy generation (calligan; zigan; calliffusion; zi2zi), yielding impressive outcomes.
Figure 1. (a) shows the character “bai” (which means “white” in Chinese) written in different fonts by various calligraphers. Each column represents a different font, and each row corresponds to a different calligrapher. (b) The first row shows calligraphy generated by Calliffusion, while the second row shows the ground truth. In the first column (san, regular script, Yan Zhenqing), the strokes in the ground truth are evenly spaced, but Calliffusion’s result is not. In the second column (ba, cursive script, Su Shi), Calliffusion’s output appears less stable than the ground truth from an aesthetic standpoint.
@@ 1. Introduction
ZiGAN (zigan) is a seminal work that applies GAN to the field of Chinese calligraphy generation. ZiGAN constructed several small-scale datasets, each containing calligraphy in a consistent style, and trained models separately on them. As a result, the trained models gained the ability to transform standard printed characters into calligraphy [^1] In the following text, we use “Calligraphy” to refer to a single character image generated by the calligraphy generation model. with the style of the specific dataset used during training. However, it cannot generate calligraphy guided by specific calligraphers, fonts, and characters. Calliffusion (calliffusion) is the first Chinese calligraphy generation model based on diffusion model(DDPM). Calliffusion uses descriptive text involving “character, font, and calligrapher” to effectively guide the generation process. However, it relies on a BERT(BERT) model pre-trained on Chinese language dataset as the text encoder, which struggles to accurately understand domain-specific terms in calligraphy, such as the names of calligraphers. For example, a calligrapher’s name might be split into two tokens. Additionally, this approach introduces extra computational cost, which is unnecessary for calligraphy generation tasks. To better accomplish this task, we proposed a multilabel mechanism, where independent classification labels correspond to the calligrapher, font, and character, and these labels are combined to control the generation process. In this way, Moyun can not only reduce the computational cost, but also control the generated calligraphy more accurately.
Calliffusion uses an Diffusion architecture based on Unet(unet), but Unet does not adequately fit the structural relationships between the strokes of the characters, which is shown in Figure 1b. The convolution operation of UNet is limited by the size of the receptive field, making it difficult to model the complex topological relationships between calligraphy strokes (such as the flying white connection and the blank space in the cursive script). Our model is based on the Vision Mamba(visionmamba), which processes images through patchify. Additionally, we incorporated the more efficient Mamba model(mamba2). Experiments demonstrate that our model provides a better fit to the structure between strokes. Vision Mamba implements long-range dependency modeling of pixel sequences through the state-space model (SSM), and its hardware-aware algorithm design is particularly suitable for processing brushstroke associations across multiple scales in calligraphy images.
The ZiGAN’s dataset(zigan) is relatively small, containing only 9 sub-datasets, with each sub-dataset consisting of 6,000 images of a single style. Moreover, this calligraphy dataset is not currently open to the public. In contrast, we collected 1.9 million single-character calligraphy images with diverse styles and detailed annotations. However, these calligraphy works come from different sources and contain a lot of noise. For example, cracks on the inscriptions, broken paper, extra ink spots, and even incomplete strokes caused by dry brush strokes. Due to varying collection conditions and different noise distributions, traditional binarization methods performed poorly. To address this, we designed a new binarization method based on SAM(SAM), which achieved better results. This method uses the prompt mechanism of Segment Anything Model to guide the segmentation of calligraphy subjects through interactive annotation, solving the problem of binarization under complex backgrounds.
@@
2. Related Work
Generative Adversarial Networks Generative Adversarial Networks (GANs) (GAN) represent a powerful framework in unsupervised learning, consisting of two competing neural networks: a generator and a discriminator. The generator aims to synthesize data that closely resembles real data by mapping from a random noise vector (typically drawn from a Gaussian or uniform distribution) to the data space. Conversely, the discriminator functions as a binary classifier, tasked with distinguishing between real data samples from the training set and synthetic data produced by the generator. The two networks are trained in an adversarial manner, engaging in a minimax game where the generator seeks to maximize the probability of the discriminator making an incorrect classification, while the discriminator strives to improve its accuracy. This dynamic leads to a progressive enhancement in the quality of generated samples as training proceeds. Since its introduction in 2014, GAN has made breakthrough progress in the field of image generation. In 2015, DCGAN (DCGAN) achieved stable training of deep convolutional GAN for the first time through a fully convolutional network architecture and a specific normalization strategy. At the application level, GAN has gone beyond the scope of image generation and has spawned diverse variants such as image-to-image translation (Pix2Pix (pix2pix)), unsupervised domain transfer (CycleGAN (CycleGAN)), and high-resolution face synthesis (StyleGAN (StyleGAN2; StyleGAN)). Research on GAN-based calligraphy generation has been ongoing for a long time, with the earliest work being zi2zi (zi2zi), an open-source project based on pix2pix (pix2pix). After that, there were calliGAN (calligan), ZiGAN(zigan), and end-to-end(zhou_end--end_2021) models. The most recent work is a style transfer model called CCST-GAN(ccstgan).
Diffusion Model The Diffusion model has emerged as one of the most widely studied and influential frameworks in image generation due to its strong generative capabilities and theoretical elegance. The foundational structure of modern diffusion models was established by the Denoising Diffusion Probabilistic Models (DDPM) (DDPM), which formalized the process of gradually adding Gaussian noise to data (forward diffusion) and then learning to reverse this process to generate samples (reverse diffusion). This paradigm demonstrated high sample quality and training stability compared to earlier generative models such as GANs. Subsequently, the Denoising Diffusion Implicit Models (DDIM) (DDIM) improved upon DDPM by introducing a non-Markovian reverse process, enabling significantly faster sampling without retraining, thus enhancing the practicality of diffusion models. Further advancements were achieved with Latent Diffusion Models (LDM) (LDM), which introduced a Variational Autoencoder (VAE) (VAE) to learn a compressed latent space representation of images. By performing the diffusion process in this lower-dimensional latent space rather than in pixel space, LDM drastically reduced computational cost and memory usage while maintaining high-fidelity generation, making it feasible to scale diffusion models to high-resolution images and diverse domains such as text-to-image synthesis. A major architectural innovation came with the introduction of Diffusion Transformers (DiT) (DiT), which replaced the traditional U-Net backbone commonly used in diffusion models with a Vision Transformer (ViT) (transformer)-based architecture. DiT demonstrated that Transformers, with their global attention mechanisms, can serve as powerful backbones for generative modeling, achieving competitive or superior performance compared to convolutional counterparts and highlighting the scalability of diffusion models with increased model capacity. Despite the rapid progress in general image generation, research on diffusion-based Chinese calligraphy generation remains relatively nascent. Early efforts in calligraphy synthesis primarily relied on GANs or rule-based systems, but recent works have begun to explore the potential of diffusion models to capture the delicate stroke structures, stylistic variations, and artistic nuances inherent in calligraphic art. The application of diffusion models in this domain presents unique challenges, including the need for fine-grained control over stroke dynamics and preservation of cultural authenticity, but also offers promising opportunities for digital heritage preservation and creative design. DP-Font(dpfont) is a Chinese font generation method based on diffusion model and physical information neural network (PINN). It improves the realism and structural accuracy of the generated fonts by introducing stroke order constraints and physical equations to guide the learning process. Compared with existing methods, DP-Font shows better performance in generating high-quality and personalized Chinese fonts. However, this work focuses on font generation rather than calligraphy generation, and the font generation task cannot focus on the style of the calligrapher. The earliest work on calligraphy is Calliffusion(calliffusion), which pioneered the application of diffusion models to Chinese calligraphy generation by employing a UNet-based architecture. Calliffusion is capable of generating calligraphy in five major fonts, and can effectively mimic the unique styles of famous calligraphers. To enable controllable generation, the model incorporates a text encoder that processes a fixed-format textual prompt describing the desired character, script, and style. These conditions are then integrated into the diffusion process through cross-attention mechanisms, allowing precise control over the generated output. Compared with Callifusion, CalliffusionV2(calliffusionV2) further introduces reference images to achieve stroke-level detail adjustment. The system adds multi-modal control capabilities and uses LoRA fine-tuning technology to quickly adapt to new styles with only 5 samples, more accurately restoring the natural strokes of complex calligraphy.
3. Methodology
@@
Figure 2 shows the architecture of “Moyun”. “Moyun” is a diffusion model based on Vision Mamba(visionmamba), optimized with Mamba2(mamba2). Furthermore, we control the generation process by the TripleLabel mechanism.
@@
Figure 2. “Moyun” architecture. The input latent noise is patched. The label is a combination of the calligrapher, font, and character. We used Mamba2-Replacement-Vision Mamba to process the patches.
3.1. Preliminaries
@@
Moyun is a diffusion model, so our training and inference follow the diffusion (DDPM; DDIM) approach. The forward process gradually adds noise to the real data, specifically to our calligraphy image $x_{0}$. The diffusion process follows the equation as shown in (1) .
(1) $q(x_{t}|x_{0})=\mathcal{N}\left(x_{t};\bar{\alpha}_{t}x_{0},(1-\bar{\alpha}_{t})\mathbf{I}\right)$
@@
(2) $x_{t}=\bar{\alpha}_{t}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\epsilon_{t}$
The diffusion model trains the reverse denoising process, and the denoising equations are as follows:
(3) $p_{\theta}(x_{t-1}\mid x_{t})=\mathcal{N}(\mu_{\theta}(x_{t}),\Sigma_{\theta}(x_{t}))$
@@
(4) $\mathcal{L}_{\text{simple}}(\theta)=\|\epsilon_{\theta}(x_{t})-\epsilon_{t}\|^{2}$
3.2. Model and Block Design
[…]
Encoder The input image $\mathbf{X}\in\mathbb{R}^{a\times a\times C}$ (where $C$ is the channel dimension) is first encoded into a compact latent representation through a pre-trained variational autoencoder (VAE)(VAE). The encoder network $\mathcal{E}$ maps the image to a latent space:
$\mathbf{Z}=\mathcal{E}(\mathbf{X})\in\mathbb{R}^{b\times b\times D}$
where $b\times b$ is the spatial resolution of the latent space, and $D$ is the channel dimension.
Patchify Module The latent tensor $\mathbf{Z}$ undergoes patchification to convert spatial information into a sequence of tokens:
$\{\mathbf{z}_{i}\}_{i=1}^{n}=\text{Patchify}(\mathbf{Z})\in\mathbb{R}^{n\times d}$
where each patch $\mathbf{z}_{i}\in\mathbb{R}^{p\times p\times D}$ is flattened into a token $\mathbf{z}_{i}\in\mathbb{R}^{d}$. The number of patches $n=\frac{b^{2}}{p^{2}}$ is determined by the patch size $p$.
moyun Block To enhance the model’s capability in capturing hierarchical structural patterns inherent in calligraphic artworks, we propose a novel block design that integrates state-of-the-art sequence modeling architectures with diffusion-based generative frameworks. The core innovation lies in replacing the conventional U-Net backbone of diffusion models with Vision Mamba(visionmamba), while incorporating the advanced Mamba2(mamba2) architecture for optimized contextual learning. Thus, each patch corresponds to a small part of the calligraphic structure, and Mamba’s efficient contextual relational ability effectively matches the relationships between patches, thereby better the model learning the structure of the calligraphy. In total, the entire moyun Block is iterated $N$ times.
3.3. TripleLabel control
To address the unique multi-conditional requirements in Chinese calligraphy generation, which include calligrapher, font, and character, we propose the TripleLabel control mechanism - a lightweight yet effective framework enabling simultaneous control over three critical aspects: calligrapher style , font structure , and character semantics. This design addresses three key challenges in calligraphy generation: (1) combinatorial generalization across calligrapher-font-character combinations, (2) efficient condition encoding without additional encoder overhead, and (3) precise condition during diffusion.
Through the TripleLabel method, each calligrapher, font, and character is mapped to a unique class label, represented by a number, and each label is Independent from others. This allows the combination of labels to generate calligraphy that the calligrapher has not written before. For example, in the training set, there is no regular script “bai” written by calligrapher Wang Xizhi, but there is an calligraphy of “bai” written by another calligrapher. During the generation process, it is still possible to generate regular script “bai” characters in the style of Wang Xizhi by providing the combinations labels of “Wang Xizhi”, “regular script” and “bai”.
In the model, the input labels are transformed into the corresponding embedding vector through three separate trainable embedding tables.
$\mathbf{e}_{\text{calli}}$ $=\text{Embed}_{\text{calli}}(l_{\text{calli}})\in\mathbb{R}^{d_{e}}$
$\vskip 5.0pt\mathbf{e}_{\text{font}}$ $=\text{Embed}_{\text{font}}(l_{\text{font}})\in\mathbb{R}^{d_{e}}$
$\vskip 5.0pt\mathbf{e}_{\text{char}}$ $=\text{Embed}_{\text{char}}(l_{\text{char}})\in\mathbb{R}^{d_{e}}$
where $l_{*}$ denotes class labels (e.g., $l_{\text{calli}}\in\{1,2,...,K_{\text{calli}}\}$).
This disentangled design leverages the additive property of Euclidean spaces for compositional control:
$\mathbf{e}_{\text{total}}=\mathbf{e}_{\text{calli}}+\mathbf{e}_{\text{font}}+\mathbf{e}_{\text{char}}\in\mathbb{R}^{d_{e}}$
@@ 3.3. TripleLabel control
enabling zero-shot generalization to unseen combinations through linear superposition. The combined embedding $\mathbf{e}_{\text{total}}$ control the generation process via a scale-shift mechanism used in DiT(DiT). It is transformed into modulation parameters $\{\alpha,\gamma,\beta\}$ through a MLP with SiLU activation. This control mechanism significantly reduces computational cost compared to introducing a new text encoder. Furthermore, subsequent experiments demonstrated that this method of control is highly effective.
4. Experiments
4.1. Dataset
@@
Figure 3. (a) shows the directory structure of dataset “Mobao”, using the calligrapher “Fan Zhongyan” as an example. “Zu” and “Zi” only have single images, while “Bai” has multiple images. (b) demonstrates the binarization process, using the character “Zu” as an example to show the steps of selecting points, obtaining the mask, and resizing.
Dataset construct To construct the large-scale annotated dataset “Mobao”, we scraped images from professional websites, and subject them to binarization processing. A total of 1.2 million raw images were initially collected, covering six major categories of Chinese calligraphic scripts: regular script, running script, cursive script, clerical script, seal script, and seal carving. Each image in the dataset is meticulously annotated with three critical attributes: calligrapher, font, and character. The calligrapher labels include 120 prominent figures from Chinese history (e.g., Wang Xizhi, Yan Zhenqing) as well as contemporary practitioners, while font labels strictly follow the six predefined script types. The dataset is organized into a hierarchical folder structure that mirrors the intrinsic relationships between calligraphers, fonts, and characters. Specifically, the top-level directories represent calligraphers, with subdirectories under each corresponding to their associated fonts (e.g., “WangXizhiregularscript” and “WangXizhicursive”). Within these font-specific folders, character labels are encoded as subdirectory names (e.g., “WangXizhiregularyi” for the character “yi”), and all images within a character folder are sequentially numbered. This hierarchical organization not only simplifies data navigation but also inherently encodes the multi-label dependency structure required for conditional generation tasks. The complete dataset hierarchical is illustrated in Figure 3(a).
To perform binarization, we employ SAM(SAM) for image segmentation. The complete processing workflow is illustrated in Figure 3(b). This approach addresses the challenges of complex backgrounds and overlapping strokes in historical calligraphy images, which are poorly handled by traditional thresholding methods. We first apply traditional computer vision techniques to identify connected regions in the image and generate point prompts for SAM. Initially, we applied Gaussian-filtered Otsu binarization method to roughly separate the foreground from the background. Subsequently, we obtained numerous connected regions from the foreground. These connected regions generally fall into two categories: one corresponds to strokes of the calligraphy, and the other consists of noise or dirt spots. To distinguish between strokes and noise, we selected a size threshold for the connected regions. This threshold was set to 100 pixels: regions larger than 100 pixels were considered as strokes and retained, while those smaller than 100 pixels were classified as noise and removed. Within these connected regions, positive points are uniformly sampled using the k-means clustering algorithm. Meanwhile, negative points are selected from areas outside of these connected domains. The number of sampled points is determined according to the following formulas:
(5) $cnt_{pos}=\max(1,\min(20,\frac{area_{regions}\times 100}{area_{total}})$
(6) $cnt_{neg}=50$
Secondly, these sampled points are then used as prompts, together with the original calligraphy images, to guide the Segment Anything Model (SAM). The model generates a corresponding segmentation mask, which we interpret as the foreground region of the image. To further refine the result, we invert the mask and fill the background region with white pixels, thereby obtaining a clean, binarized image representation. Finally, a resizing and padding process is performed to standardize the image dimensions. Specifically, the image is first resized such that its longer side is scaled to 256 pixels while preserving the original aspect ratio. Subsequently, the resized image is centered within a $256\times 256$ white canvas by padding the remaining space with white pixels. This results in a final square image of size $256\times 256$ , suitable for downstream processing or model input requirements.
@@
dataset analyse After processing the dataset, we performed a comprehensive analysis of the resulting images. A total of 1,929,393 images were obtained, representing calligraphic works from six different fonts, created by 2,681 distinct calligraphers, and covering 4,660 unique Chinese characters. Notably, some images lacked attribution to a specific calligrapher and were therefore classified under the “anonymous” category.
@@
To further evaluate the characteristics of our dataset, we conducted a statistical analysis across three key dimensions: calligrapher, font, and character. The results confirm the large scale of our dataset. For instance, the calligrapher with the most extensive collection, “Wang Xizhi”, contributes a total of 124,854 images; the font with the largest representation, running script, accounts for 668,040 images; and the most frequently occurring character, “shu” (meaning “book” in Chinese), appears in 11,949 images. These statistics demonstrate that the dataset provides a rich and diverse source of data for models to learn both the structural and stylistic features of Chinese calligraphy.
Figure 4. (a) indicates the number of calligraphy per calligrapher.(y-axis: $\ln$ of calligraphy count) (b) indicates the number of calligraphy per character. Both the red and green lines mark the top $10\%$ and $50\%$ thresholds, respectively, highlighting the long-tail data imbalance.
However, we also observe notable imbalances in the data distribution. Nearly half of the calligraphers are associated with fewer than ten images, and approximately half of the characters are represented by less than one hundred samples. This suggests that the dataset follows a long-tail distribution as shown in Figure 4, where a small number of categories are heavily represented while the majority have limited coverage. The red and green lines in the figure denote the top $10\%$ and $50\%$ thresholds, further highlighting the skewed nature of the distribution across both calligrapher and character dimensions.
4.2. Experiment Setup
Experiment Dataset To ensure the reliability and fairness of our experiments, we constructed a balanced subset from the full dataset to serve as the experimental benchmark. The construction process is as follows: We selected 40 calligraphers who have written a common set of 40 characters. For each calligrapher, 90% of the characters were assigned to the training set, while the remaining $10\%$ characters were reserved for testing. This ensures that each test character is unseen during training for that specific calligrapher. However, other characters written by the same calligrapher, as well as the same characters written by other calligraphers, are still included in the training set. This design enables the model to learn both the structural variations of characters and the stylistic characteristics of individual calligraphers.
Given the varying number of available images per character across calligraphers, we randomly sampled up to four images for each character-calligrapher pair. In total, we obtained 12,985 images, with 11,689 allocated to the training set and 1,296 to the test set. This careful curation effectively mitigates the negative impact of the long-tail distribution, resulting in a more balanced and representative experimental dataset. Model Specifics For the choice of the variational autoencoder (VAE)(VAE), we employed the same pre-trained VAE as used in Latent Diffusion Models (LDM)(LDM). Specifically, the input image of size $256\times 256\times 3$ were first encoded into a lower-dimensional latent space of size $32\times 32\times 4$.